List all the possible samples and the corresponding mean

1
Expert's answer
2022-06-06T14:40:05-0400

We have population values 1,2,3,4,5, population size N=5 and sample size n=2.

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5}{5}=3"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{4+1+0+1+4}{5}=2"



Select a random sample of size 2 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_2=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,2 & 3/2 \\
 \hdashline
 2 & 1,3 & 4/2 \\
 \hdashline
 3 & 1,4 & 5/2 \\
 \hdashline
 4 & 1,5 & 6/2 \\
 \hdashline
 5 & 2,3 & 5/2 \\
 \hdashline
 6 & 2,4 & 6/2 \\
 \hdashline
 7 & 2,5 & 7/2 \\
 \hdashline
 8 & 3,4 & 7/2 \\
 \hdashline
 9 & 3,5 & 8/2 \\
 \hdashline
 10 & 4,5 & 9/2 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 3/2 & 1/10 & 3/20 & 9/40 \\
 \hdashline
 4/2 & 1/10 & 4/20 & 16/40 \\
 \hdashline
 5/2 & 2/10 & 10/20 & 50/40 \\
 \hdashline
 6/2 & 2/10 & 12/20 & 72/40 \\
 \hdashline
 7/2 & 2/10 & 14/20 & 98/40 \\
 \hdashline
 8/2 & 1/10 & 8/20 & 64/40 \\
 \hdashline
 9/2 & 1/10 & 9/20 & 81/40 \\
 \hdashline
\end{array}"


Mean of sampling distribution 


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{60}{20}=3=\mu"


The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{390}{40}-(3)^2=\dfrac{3}{4}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"